OpenStudy (loser66):

You are given the pdf \(\phi(x)=\dfrac{1}{4\sqrt{2\pi}}e^{-x^2/32}\) \(\-\infty

OpenStudy (anonymous):

I'm not exactly sure about the details of deriving the CDF of the normal distribution given the PDF (as shown on the wiki page, http://en.wikipedia.org/wiki/Normal_distribution ; I'm not terribly familiar with the error function \(\text{erf}\)). You can verify, though, that \(\Phi'(x)=\phi(x)\) by the fundamental theorem of calculus. For #2, you have \[\begin{align*} G(w)&=P(W\le w)\\ &=P(X^2\le w)\\ &=P(|X|\le \sqrt w)\\ &=P(-\sqrt w\le X\le \sqrt w)\\ &=P(X\le \sqrt w)-P(X\le -\sqrt w)\\ &=\Phi(\sqrt w)-\Phi(-\sqrt w) \end{align*}\] Again, this assumes you have the closed form of \(\Phi\), which is the main difficulty here (in my opinion). #3 and #4 should be easy enough once you determine \(G(w)\).

OpenStudy (loser66):

Yes, the most difficult part is part 1 :(

OpenStudy (anonymous):

Let's see what we can do. Here's the definition for the error function: http://mathworld.wolfram.com/Erf.html \[\newcommand{\erf}{\text{erf}} \erf(x)=\frac{2}{\sqrt \pi}\int_0^x e^{-t^2}dt\] Here's what we want to find: \[\begin{align*} \Phi(x)&=\int_{-\infty}^x\frac{1}{4\sqrt{2\pi}}e^{-t^2/32}~dt\end{align*}\] Set \(\dfrac{t}{4\sqrt 2}=s\), then \[\begin{align*} \Phi(x)&=\frac{1}{\sqrt\pi}\int_{-\infty}^{x/4\sqrt2} e^{-s^2}~ds\\\\ &=\frac{1}{2}\left(\underbrace{\frac{2}{\sqrt\pi}\int_{0}^{x/4\sqrt2} e^{-s^2}~ds}_{\text{def. of erf}}\right)+\underbrace{\frac{1}{\sqrt\pi}\int_{-\infty}^0 e^{-s^2}~ds}_{\text{this is a known result}}\\\\ &=\frac{1}{2}\erf\left(\frac{x}{4\sqrt2}\right)+\frac{1}{2\sqrt\pi}\int_{-\infty}^\infty e^{-s^2}~ds \end{align*}\]

OpenStudy (loser66):

Wow, it is a mess. :) Thanks for the link. :)

OpenStudy (loser66):

I know \(\int_{0}^{\infty}e^{-x^2}dx=\dfrac{\sqrt{\pi}}{2}\)

OpenStudy (loser66):

That is Gaussian integral.

OpenStudy (anonymous):

Oops, I meant \[\int_{-\infty}^\infty e^{-x^2}dx=\sqrt{\pi}\] But okay, that's good. Less work!

OpenStudy (anonymous):

Alright, so now to find \(G(w)\) you have \[\begin{align*} G(w)&=\Phi(\sqrt w)-\Phi(-\sqrt w)\\\\ &=\left[\frac{1}{2}\erf\left(\frac{\sqrt w}{4\sqrt2}\right)+\frac{1}{2}\right]-\left[\frac{1}{2}\erf\left(\frac{-\sqrt w}{4\sqrt2}\right)+\frac{1}{2}\right]\\\\ &=\frac{1}{2}\erf\left(\frac{\sqrt w}{4\sqrt2}\right)+\frac{1}{2}\erf\left(\frac{\sqrt w}{4\sqrt2}\right)\\\\ &=\erf\left(\frac{\sqrt w}{4\sqrt2}\right) \end{align*}\] The second step is due to \(\erf(x)\) being an odd function. From here you can find \(g(w)\) by taking the derivative. Recall that \[\begin{align*}G(w)&=\erf\left(\frac{\sqrt w}{4\sqrt2}\right)\\\\ &=\frac{2}{\sqrt\pi}\int_0^{\sqrt w/4\sqrt2} e^{-t^2}~dt\\\\ G'(w)&=\frac{2}{\sqrt\pi}e^{-(\sqrt w/4\sqrt2)^2}\times\frac{d}{dw}\left[\frac{\sqrt w}{4\sqrt2}\right]\\\\ g(w)&=\frac{1}{4\sqrt{2\pi}}\frac{e^{-w/32}}{\sqrt w} \end{align*}\]

OpenStudy (loser66):

Wow!! It's a big help. :)

OpenStudy (anonymous):

Showing that \(\displaystyle\int_0^\infty g(w)~dw=1\) is simple enough, I think. It'll involve the error function. Quick tip: substitute \(y=\sqrt w\).

OpenStudy (loser66):

Yes, Sir. I know how to do it. :)